how many ounces of a 45% alcohol solution must be mixed with 16 ounces of a 50% alcohol solution to make a 49% alochol solution

Answer by mananth(16946)

-------------- percent ---------------- Amount
Alcohol I 50 ---------------- 16 ounces
Alcohol II 45 ---------------- x ounces
Mixture 49 ---------------- 16 + x ounces

50 * 16 + 45 x = 49 ( 16 + x )
800 + 45 x = 784 + 49 x
45 x -49 x = -800 + 784
-4 x = -16
/ -4
x = 4 ounces Alcohol II

m.ananth@hotmail.ca